OpenStudy (anonymous):

A playground is surrounded by a fence 2.0 m high. A boy who plays in the playground kicks a ball vertically upward from a height of 0.2 m with a speed of 10 m/s. How much time does the ball spend above the height of the fence?

OpenStudy (anonymous):

do you know the equation for height in meters? i forget. somethng like \[h(t)=-4.9t^2+v_0t+h_0\] right?

OpenStudy (mertsj):

Yes. That is right, smartie.

OpenStudy (anonymous):

ok then if that is it solve \[-4.9t^2+10t+.2>2\]

OpenStudy (anonymous):

ick this is a real pain

OpenStudy (anonymous):

\[-4.9t^2+10t-1.8=0\] \[49t^2-100t+18=0\] difference of the roots is \[\frac{\sqrt{b^2-4ac}}{a}\]

OpenStudy (mertsj):

I think it might be easier to set it equal to 0 and solve the equation. You will get 1.84 and .2 The ball will be higher when the x value is between those two values. So 1.84-.2=1.64 sec it will be higher than the fence.

OpenStudy (anonymous):

i get \[\frac{2 \sqrt{1618}}{49}\]

OpenStudy (anonymous):

1.64 seconds rounded

OpenStudy (anonymous):

my answer 0.82s. is it correct?

OpenStudy (anonymous):

no , i think it should be 1.64

OpenStudy (anonymous):

ok2... my fault..sory..thanx all my friends

OpenStudy (anonymous):

yw